OpenStudy (anonymous):

ln5^x-ln5^(x-1) plsssssss help in dire need

OpenStudy (anonymous):

instructions are to simplify completely

OpenStudy (danjs):

Do you recall the log properties for how you can move things around?

OpenStudy (anonymous):

nope

OpenStudy (danjs):

i would do this first \[\large \ln(b^a)=a*\ln(b)\]

OpenStudy (anonymous):

got it

OpenStudy (danjs):

\[x*\ln(5) - (x-1)*\ln(5)\] factor out the ln(5) and simplify, what that gie ya

OpenStudy (anonymous):

ln5(-1)

OpenStudy (anonymous):

i meant 1

OpenStudy (danjs):

yeah good , remember the basic log rules and you are good, first 3 here http://www.purplemath.com/modules/logrules.htm

OpenStudy (anonymous):

thank u friend!!!